Explain how you can find the product of 40x16 by breaking apart the numbers
Found 2 solutions by checkley77, drk:
Answer by checkley77(12844)

40(10+6)=400+240=640 ans.

Answer by drk(1908)

There are several ways to solve this.I will show the Russian duplication method also called Russian peasant multiplication.
40 x 16
double the first and 1/2 the second
80 x 8
double the first and 1/2 the second
160 x 4
double the first and 1/2 the second
320 x 2
double the first and 1/2 the second
640 x 1 = 640.
--
If you meant something like (40)*(10+5+1)
then distribute the 40 to get
400 + 200 + 40
which equals 640.